OpenStudy (anonymous):

Jason considered two similar televisions at a local electronics store. The generic version was based off the brand name and was 3/4 the size of the brand name. If the generic television set is 12 inches by 24 inches, what are the dimensions of the brand name television?

OpenStudy (mysesshou):

12*(4/3) by 24*(4/3) =16 by 32 imagine: 12/x = 3/4 and solve for x. 24/x = 3/4 solve for x. to check: 3/4 of 16 is 12 because 16 divided by 4 times 3 is 12.